OpenStudy (anonymous):

Which of the following shows the equation below correctly solved for x? Plz help ASAP And explain how u got answer Question once u open

OpenStudy (anonymous):

6xy=7x+3y+z

OpenStudy (anonymous):

X= 7y+z/ 6y-0. This is one choice X=7y+z/ 6y+3

OpenStudy (anonymous):

X= 3y+z/ 6y-7 none of the above

OpenStudy (whpalmer4):

\[6xy=7x+3y+z\]Subtract \(7x\) from both sides to get all terms involving \(x\) on left hand side \[6xy-7x=3y+z\]Now factor out \(x\) on left side \[x(6y-7)=3y+z\]Solve for \(x\) by dividing both sides by \(6y-7\) \[x=\frac{3y+z}{6y-7}\]

OpenStudy (whpalmer4):

Your last choice is an improperly written version of this. To make it correct, you would need to use the equation editor button, or parentheses to show the correct order of evaluation: x = (3y+z)/(6y-7). Otherwise, what you wrote is actually \[x=3y+\frac{z}{6}y-7\] because multiplication and division take precedence over addition and subtraction (remember PEMDAS?)

OpenStudy (anonymous):

Which of the following shows the equation below correctly solved for g? f =1/2 g − 7h